OpenStudy (anonymous):

Solve the following exponential equation. Please show all of your work. 2^(4x-7)=128

OpenStudy (rulnick):

4x-7 = log_2 128 4x - 7 = 7 4x = 14 x = 14/4 x= 7/2 done :)

OpenStudy (anonymous):

Thanx!

OpenStudy (rulnick):

welcome.

OpenStudy (anonymous):

2^(4x-7)= 2^7 ->4x -7 = 7 => x = 14/4 = 7/2